OpenStudy (anonymous):

At 3 p.m. Coretta's shadow is 1.05 meters long. Her height is 1.68 meters. At the same time, a tree's shadow is 6.58 meters long. How tall is the tree? Round your answer to the nearest hundredth. I think its 4.11 meters but can someone please check it for me?

OpenStudy (nikvist):

\[\frac{C_S}{C_H}=\frac{T_S}{T_H}\quad\Rightarrow\quad T_H=T_S\frac{C_H}{C_S}=6.58\cdot\frac{1.68}{1.05}=10.53\]

OpenStudy (anonymous):

1.05/1.68=6.58/x, 1.05x=1.68*6.58, x=10.53